OpenStudy (anonymous):

find a vector equation and parametric equations through the point (1,0,6) and perpendicular to the plane x + 3y + z =5

OpenStudy (amistre64):

is the point on a line? thru the plane?

OpenStudy (anonymous):

ummm what?

OpenStudy (amistre64):

if so, the normal of the plance is the perp to it; strip off your coeefs for the line vector: <1,3,1> would be the vector that is perpenducular to the plane

OpenStudy (amistre64):

since 1,0,6 is not on the plane im going to assume its on the line perp to it (the perp line is called a normal)

OpenStudy (anonymous):

okay I'm with you so far

OpenStudy (amistre64):

x = Px + 1t y = Py + 3t z = Pz + 1t given the point (1,0,6) this makes x = 1 + 1t y = 0 + 3t z = 6+ 1t

OpenStudy (anonymous):

okay but why are the points (1,3,1) perpendicular to the plane?

OpenStudy (amistre64):

its a result of how the equation of a plane is defined

OpenStudy (amistre64):

the perp to the plane is a vector that dot producted to every line in the plane

OpenStudy (amistre64):

so given a point (Px,Py,Pz) in space: we can construct any vector from it by taking any random point (x,y,z) and subtracting the 2 of them: (x-Px, y-Py, z-Pz) is every vector that comes out from our point into space

OpenStudy (amistre64):

to elimante all the vectors that are not in a plane; we need to establish a perp vector to it so that when we dot it to these vectors we get 0, so all vectors of the form: (x-Px,y-Py,z-Pz) dot (Nx,Ny,Nz) = 0 Nx(x-Px) +Ny(y-Py) +Nz(z-Pz) = 0 is the equation of the plane

OpenStudy (anonymous):

okay makes sense!

OpenStudy (amistre64):

good :)

OpenStudy (anonymous):

you're way better at explaining things than my book...

OpenStudy (amistre64):

thats becasue im crazy lol

OpenStudy (anonymous):

lol, does normal mean perpendicular?

OpenStudy (mathmate):

yes.

OpenStudy (mathmate):

Actually, yes in this context. The word normal is used for other things in maths elsewhere.